Circular No. 6555

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) E. Kreysa, W. Altenhoff, and C. G. T. Haslam, Max-Planck-Institut

für Radioastronomie, Bonn; and A. Sievers, Institut de Radio Astronomie Millimétrique (IRAM), Granada, report the detection of comet C/1995 O1 in the continuum at 250 GHz with the IRAM 30-m radio telescope on Pico Veleta. The preliminary results (with internal/absolute errors) are: Feb. 1.355 UT, 102 ± 6/15.3 mJy; 2.345, 92 ± 2/13.8 mJy. Multiple scans through the comet give no indication for a significant beam broadening. Under the assumption of a nuclear diameter of order 30 km, more than 90 percent of the observed signal must come from the (ice) particle halo.”

C. W. Hergenrother and S. M. Larson, Lunar and Planetary Laboratory; and B. G. Marsden, Center for Astrophysics, report the appearance of changing structure in jets of C/1995 O1 from processed R-band CCD images obtained with the 1.2-m Whipple Observatory telescope on Feb. 1.5 UT. Confirming CCD observations were obtained by R. Tucker (Goodricke-Pigott Observatory) with a 0.36-m telescope on Feb. 1.5 and 3.5. Bright linear jets are seen at p.a. 25^{o}, 166^{o}, and 208^{o} and fainter ones at 91^{o} and 306^{o}. An additional bright but curved jet is seen close to and south of the central condensation on both days with different orientations — though apparently related to the two southern jets, which contain at least five periodic brightness enhancements 8′′ apart extending along each jet. Each enhancement in both of these jets has curved tail-like extensions fading towards the approximate anti-solar direction. Further observations providing better temporal coverage are encouraged to see if the production mechanism is rotationally modulated, or if some other process is responsible.

H. Mikuž, Crni Vrh Observatory, reports the following total integrated magnitudes from V -filtered CCD images taken through camera lenses of focal lengths 180 (f/2.8) and 90 mm (f/4): Jan. 16.22 UT, 2.8 (180); 17.20, 2.8 (180); 29.20, 2.2 (90); 31.21, 2.3 (90); Feb. 3.20, 2.0 (90, 180); 4.21, 1.9 (90, 180). The coma diameters were ∼ 50′ in the 180-mm lens and ∼ 60′ in the 90-mm lens. Further naked-eye m_{1} estimates: Jan. 9.53 UT, 3.1 (A. Hale, Cloudcroft, NM); 13.21, 2.8 (R. Haver, Frasso Sabino, Italy); 17.45, 2.7 (J. E. Bortle, Stormville, NY); 21.50, 2.5 (G. W. Kronk, Troy, IL); 26.21, 2.4 (M. Plsek, Lelekovice, Czech Rep.); 27.23, 2.5 (R. J. Bouma, Groningen, The Netherlands); 31.24, 2.3 (B. H. Granslo, Fjellhamar, Norway); Feb. 2.17, 2.0 (K. Sárneczky, Budapest, Hungary; 7^{o} tail in p.a. 330^{o}); 3.21, 1.9 (V. Znojil, Brno, Czech Rep.); 4.14, 1.8 (A. Baransky, Kiev, Ukraine).

1997 February 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green